# Machine learning analysis of anomalous diffusion

### v1

@@ 1 Introduction
where the diffusion exponent $\alpha$ categorizes the diffusion as either superdiffusive ($\alpha>1$) or subdiffusive ($\alpha<1$). Another distinct aspect involves the stochastic process in anomalous diffusion, which contrast significantly with the Wiener process associated with Brownian motion. For this reason, anomalous diffusion trajectories often exhibit long-range correlations and memory effects, or demonstrates the “anomalous yet Brownian" diffusion characteristic [35, 36, 37, 38, 39, 40]. As a response, numerous theoretical models used to describe the evolution of anomalous diffusion have been proposed, such as the continuous-time random walk (CTRW)[41, 42] and fractional Brownian motion (FBM)[43]. CTRW stands out for its integration of random steps with variable waiting times, effectively modeling memory and non-Markovian effects. FBM extends traditional Brownian motion by incorporating a Hurst exponent, allowing it to describe random processes exhibiting persistent or anti-persistent characteristics. A third key distinction between anomalous diffusion and Brownian motion arises in complex environments, where a random walker may exhibit varying diffusion states over time due to intrinsic or extrinsic changes, leading to heterogeneous diffusion dynamics[44, 45, 46, 13, 14, 47, 48, 49, 50, 51, 52, 53, 54, 55]. For instance, the Hfq protein in Escherichia coli cells can continuously transition among three diffusion states with different diffusion coefficients and mean dwell times [45]. Similarly, in the microtubule network of DU145 cells [44], the transports of endosomes can vary among Lévy walk (LW) [56] and Brownian motion. These complexities emphasize the need for advanced methods to decode the diffusion data and unravel the diverse nature of anomalous diffusion.
@@ 2.1 Inference of diffusion parameters
Therefore, before the utilization of machine learning, numerous methods have been developed to infer these diffusion parameters. A classic example is the use of time-averaged mean squared displacement (TA-MSD) to determine the diffusion exponent $\alpha$ based on Eq. (1). Another instance is the analysis of power spectral density (PSD), which provides the frequency-dependent information of the trajectories [57, 58, 30, 59]. In addition, velocity autocorrelation function (VACF) [60], fractional integral moving average (FIMA) process [61, 62, 63] and $p$-variation test[64] are also extensively employed when calculating the diffusion parameters. These traditional methods offer straightforward and highly interpretable analytical approaches that help us uncover the physical information embedded in diffusion trajectories.
@@
To address this challenge, machine learning techniques have been introduced to the inference of diffusion parameters due to its powerful ability to handle complex data in recent years. In particular, inferring diffusion parameters essentially constitutes a regression or classification problem, fitting well within the standard supervised learning paradigm. For clarity in our discussion, we have organized these ML approaches into two primary categories in this section: classical machine learning and deep learning. Within classical machine learning, we highlight significant contributions using techniques such as feature-based models[109, 110, 111, 112] and Bayesian methods[113, 114, 115, 116, 117, 118]. Regarding deep learning, we focus on innovative applications that employ deep neural networks, including convolutional neural networks [119, 120, 121, 122, 123, 124, 125, 126, 127], recurrent neural networks[128, 129, 130, 131, 132], graph neural networks[133, 134, 135], to analyze diffusion trajectories and infer the parameters.
@@ 2.1.1 Classical machine learning
Classical machine learning encompasses a range of algorithms and techniques that predate the extensive use of deep neural networks. These methods are grounded in established statistical learning theories and are characterized by their simplicity, interpretability, and lower computational requirements. For the analysis of anomalous diffusion, classical machine learning primarily relies on feature-based models or Bayesian approaches to determine diffusion parameters.
As highlighted in figure 2(a), feature-based models operate on manually extracted features from diffusion trajectories and utilize classic machine learning algorithms such as decision trees and multilayer perceptrons for parameter inference. By harnessing handcrafted features that capture the essential characteristics of the diffusion process, such as fractal dimension, maximal excursion, and mean gaussianity [70, 71, 72], these models provide a structured way to analyze and interpret diffusion data. Typical related works include the following:
Figure 2: (a) Schematic flowchart of the feature-based methods: a set of features is extracted from raw trajectories and used as the input for feature-based machine learning models, from [73]; (b) A workflow of random forest algorithm for the single trajectory characterization tasks, from[109]; (c) Schematic diagram of the Markovian approximation of discrete-time Lévy walks (top); comparisons between Bayesian inference and MSD fitting across different signal-to-noise ratio conditions when predicting diffusion exponents (bottom), from [115].
  • A study by G. Muñoz-Gil et al. [109] employs the random forest (RF) algorithm for the single trajectory characterization tasks, which is a classical machine learning method by aggregating results of multiple decision trees through voting or averaging As shown in figure 2(b), upon the supervision of labeled simulated trajectories from various diffusion models, the RF model can be trained to infer diffusion exponent and theoretical model after preprocessing and feature extraction of these trajectories. The trained model is then applied to experimental data, including the movement of mRNA molecules in bacterial cells and membrane receptor diffusion in living cells. The results show that the RF model achieves performance superior to traditional methods, particularly in the analysis of nonergodic diffusion processes.
@@
  • Another application of classical machine learning is the use of extreme learning machine (ELM) in C. Manzo’s work [111]. This method is tailored for single-layer feedforward networks, simplifying the learning process by randomly assigning weights between the input and hidden layers. In this study, ELM is integrated with feature engineering to analyze anomalous diffusion trajectories. The input to the ELM model includes features such as time-normalized mean absolute displacement, mean square displacement, and cumulative square displacement. These features are efficiently processed by ELM to infer critical diffusion parameters like diffusion exponent and diffusion model. The model performance underscores the effectiveness of ELM in handling complex, feature-rich datasets, demonstrating its potential to deliver rapid and accurate analysis of diffusion behaviors.
@@
In practice, the effectiveness of these feature-based models heavily depends on the statistical features extracted from trajectory data, such as TA-MSD, displacement moments, or other engineered features that characterize the diffusion dynamics, which will be discussed in detail in section 3.1. As these features typically possess strong statistical or physical significance, they provide feature-based methods with enhanced interpretability compared to deep learning. However, selecting the optimal combination of features is challenging, which requires deep domain expertise and extensive experimental analysis to identify the most significant characteristics of diffusion properties. In this regard, H. Loch-Olszewska et al. underscores the impact of feature choice on machine learning classification of anomalous diffusion [112]. The study shows that while combining multiple feature sets can enhance the performance of classifier, no single feature set can universally excel across all diffusion models. This variability highlights the profound impact of feature selection, and also suggests that feature-based machine learning methods may have limitations in their universality for the analysis of anomalous diffusion.
Beyond the feature-based methods, Bayesian inference represents another powerful classic machine learning strategy [113, 114, 115, 116, 117, 118], renowned for its robust interpretability and analytical depth in parameter inference for anomalous diffusion. Bayesian methods offer a probabilistic perspective that enables the integration of prior knowledge with observed data, facilitating a dynamic update of our understanding as new data becomes available. This probabilistic framework employs posterior distributions derived from Bayesian updates to quantify the uncertainty in parameter estimates effectively. By updating the prior beliefs with likelihoods calculated from new data, Bayesian inference continually refines the estimates of model parameters. This approach not only aids in selecting the model that best explains the data but also enhances the understanding of the uncertainty and variability inherent in the model predictions.
One representative work of Bayesian inference for anomalous diffusion is conducted by S. Park et al. [115], which enables an efficient parameter estimation of Lévy walks approximated by the hidden Markov model [figure 2(c)]. The Bayesian approach in this work allows for the inference of the power-law exponent of flight-time distribution and the classification of different diffusion models. This method excels in handling noise and incomplete data, providing robust estimations even when traditional methods like MSD fitting fail. Another work by J. Krog et al. [117] combines Bayesian inference with an information content model check to analyze stochastic processes described by Langevin equations. The authors apply this approach for SPT data to estimate model parameters and compare different candidate models. The results provide not only the most probable parameters but also their uncertainties. In addition, H. Seckler et al. have combined the Bayesian framework with deep learning for error estimation in diffusion trajectory analysis [114]. These studies demonstrate that Bayesian inference provides a flexible and powerful paradigm capable of integrating various tools to achieve the analysis of anomalous diffusion.
@@ 2.1.3 Further discussions
While both classic machine learning and deep learning have demonstrated outstanding performance in the inference of diffusion parameters, there are critical considerations that need to be addressed for these methods to reach their full potential. First, the application of machine learning in this field primarily falls within the paradigm of supervised learning, where the task is either a classification or regression problem. Therefore, large and high-quality datasets with accurate labels are required when training these models. However, obtaining labels from experimental data can be challenging. To address this issue, the main current approach is the simulation-based method [109, 110, 111, 112, 114, 120, 121, 122, 119, 123, 126, 124, 125, 128, 129, 130, 131, 132, 133, 135, 134, 136]. This method uses various known mathematical diffusion models, such as CTRW, FBM, and LW, to simulate trajectories. These trajectories can be labeled based on the diffusion models and corresponding simulation parameters. After training the machine learning models on simulated trajectories, these models can be subsequently applied to the analysis of real experimental data. Although this method has shown excellent performance so far, it faces an inherent challenge due to the natural differences between simulated and experimental data. On one hand, experimental trajectories often exhibit higher levels of noise, stochasticity, and complexity compared to their simulated counterparts. On the other hand, many characteristics of experimental trajectories cannot be fully captured and simulated by known diffusion models. This clearly limits the generalizability of machine learning models trained solely on synthetic data. Thus, the improvement and optimization of the simulation-based method will be a critical issue for future research.
@@
    The interpretability of models represents a crucial distinction between classical machine learning and deep learning approaches. Classical techniques like decision trees and random forests offer higher transparency due to their reliance on explicitly defined features. As stated before, these crafted features are selected with substantial domain knowledge. They often root in well-understood physical or statistical properties of the diffusion process, such as TA-MSD and VACF. This level of clarity is invaluable, especially in scientific research where understanding the underlying mechanisms is as crucial as the predictive performance of the model. On the contrary, deep learning models are often criticized for their “black box" nature, as their internal workings remain largely opaque. Nevertheless, it is possible to gain insights into the role of each component within deep neural networks through methods such as ablation studies [137, 138]. These techniques can help elucidate the underlying mechanisms by which neural networks infer diffusion parameters.
@@ 2.2 Segmentation of heterogeneous diffusion dynamics
As highlighted in the introduction, a significant distinction between anomalous diffusion and Brownian motion emerges in complex environments, where a random walker may exhibit varying diffusion states over time due to intrinsic or extrinsic changes, leading to heterogeneous diffusion dynamics [44, 45, 46, 13, 14, 47, 48, 49, 50, 51, 52, 53, 54, 55]. Compared to single-state trajectories with fixed properties, heterogeneous dynamics involve transitions among multiple diffusion states. This makes the accurate analysis of multi-state trajectory data much more complex and thus poses significant challenges in the single trajectory characterization. Segmentation of heterogeneous diffusion dynamics is therefore a crucial task in the analysis of anomalous diffusion. This task involves two primary subtasks:
@@ 2.2.2 Point-wise method
While the sliding window method offers a systematic approach to segmenting trajectories, it inherently suffers from limitations, particularly regarding the predefinition of window length. The choice of window length is often a trade-off: larger windows might mask rapid transitions between states, while smaller ones could increase noise and lead to misclassifications. These constraints become more pronounced when dealing with complex and heterogeneous diffusion dynamics, where transitions between states do not necessarily follow uniform lengths.
@@
Table 1: Overview of selected methods for the segmentation of anomalous diffusion.
@@
Another example of point-wise method is the U-AnDi model proposed by X. Qu et al. [138]. As illustrated in figure 6(c), this model combines dilated causal convolutions to capture long-range dependencies, gated activation units for time-series information filtering, and a U-Net architecture to achieve fine-grained segmentation performance. Similar to the STEP method, U-AnDi predicts diffusion parameters for each point along a trajectory, including the diffusion exponent and diffusion model. As shown in figure 6(d), after training on simulated datasets, U-AnDi is further tested on experimental data involving the diffusion of transmembrane proteins on cell membrane surfaces. These experimental trajectories include two diffusion models: CTRW and FBM, with U-AnDi demonstrating excellent segmentation performance for both. In addition to the STEP and U-AnDi models, other machine learning approaches like RNN [152] and GNN [135] also employ point-wise methods to segment trajectories effectively. Detailed information regarding these models is listed in Table 1.
@@ 2.3 Anomalous Diffusion Challenge
The AnDi Challenge is organized by G. Muñoz-Gil et al. and is hosted on the CodaLab platform [157]. This competition features two distinct versions. The first iteration in the year 2020 focuses primarily on the inference of diffusion parameters, including the diffusion exponent and diffusion model [29]. While this version includes a segmentation task, its simplistic design does not fully replicate the complexities encountered in real-world scenarios. Consequently, the subsequent iteration of the challenge in the year 2024 shifts its emphasis towards the segmentation of heterogeneous diffusion dynamics [158], aiming to address this complex problem more directly and effectively. Details of the AnDi Challenge are listed in Table 2, including the task descriptions, evaluation metrics, number of teams for each task, and links to related works. The ranking information can be found on the official competition website [89] and is not listed in the table here for brevity. In the following, we provide an overview of each version of the AnDi Challenge, focusing on their specific objectives and respective datasets:
@@
$2^{\rm nd}$ This challenge is organized along two tracks: Video track: For analysis of raw SPT videos. Trajectory track: For the analysis of trajectory data. Each track has two tasks: Task 1 (Single-trajectory task): For each trajectory, participants are supposed to list the changepoints. After that, they need to determine the diffusion coefficient, diffusion exponent, and motion type for each segment. Task 2 (Ensemble task): For each experimental condition, participants are required to provide the model type, number of states, mean and standard deviation of diffusion coefficients, and diffusion exponents for each state. Task 1: JSC, RMSE, MSLE, MAE, F1-Score Task 2: Wasserstein distance Video track: Task 1: 3 teams Task 2: 3 teams Trajectory track: Task 1: 18 teams Task 2: 11 teams [158]
@@
However, despite its significant contributions, the AnDi Challenge highlights a critical gap in this field: the reluctance within the scientific community to share real experimental data openly. Numerous studies on the machine learning characterization of anomalous diffusion rely heavily on the simulation-based method due to the lack of accessible real datasets. This situation constrains the development of machine learning tools that could be effectively applied to actual experimental analysis. To address this issue, there is a pressing need for the SPT research community to actively promote data sharing. Encouraging the open source of real experimental data would not only validate the applicability of these models in practical settings, but also accelerate the discovery of novel insights into the dynamics of anomalous diffusion processes.

### v2

@@ 1 Introduction
where the diffusion exponent $\alpha$ categorizes the diffusion as either superdiffusive ($\alpha>1$) or subdiffusive ($\alpha<1$). Another distinct aspect involves the stochastic process in anomalous diffusion, which contrast significantly with the Wiener process associated with Brownian motion. For this reason, anomalous diffusion trajectories often exhibit long-range correlations and memory effects, or demonstrates the “anomalous yet Brownian" diffusion characteristic [35, 36, 37, 38, 39, 40]. As a response, numerous theoretical models used to describe the evolution of anomalous diffusion have been proposed, such as the continuous-time random walk (CTRW) [41, 42] and fractional Brownian motion (FBM) [43]. CTRW stands out for its integration of random steps with variable waiting times, effectively modeling memory and non-Markovian effects. FBM extends traditional Brownian motion by incorporating a Hurst exponent, allowing it to describe random processes exhibiting persistent or anti-persistent characteristics. A third key distinction between anomalous diffusion and Brownian motion arises in complex environments, where a random walker may exhibit varying diffusion states over time due to intrinsic or extrinsic changes, leading to heterogeneous diffusion dynamics [44, 45, 46, 13, 14, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57]. For instance, the Hfq protein in Escherichia coli cells can continuously transition among three diffusion states with different diffusion coefficients and mean dwell times [45]. Similarly, in the microtubule network of DU145 cells [44], the transports of endosomes can vary among Lévy walk (LW) [58] and Brownian motion. These complexities emphasize the need for advanced methods to decode the diffusion data and unravel the diverse nature of anomalous diffusion.
@@ 2.1 Inference of diffusion parameters
Therefore, before the utilization of machine learning, numerous methods have been developed to infer these diffusion parameters. A classic example is the use of time-averaged mean squared displacement (TA-MSD) to determine the diffusion exponent $\alpha$ based on Eq. (1). Another instance is the analysis of power spectral density (PSD), which provides the frequency-dependent information of the trajectories [59, 60, 30, 61]. In addition, velocity autocorrelation function (VACF) [62], fractional integral moving average (FIMA) process [63, 64, 65] and $p$-variation test [66] are also extensively employed when calculating the diffusion parameters. Furthermore, Bayesian inference [92, 93, 94, 95, 96, 97] has emerged as another powerful method for parameter estimation in the study of anomalous diffusion. It provides a probabilistic framework where prior knowledge about the system can be updated with observed data. This allows for continuous refinement of parameter estimates and offers a measure of uncertainty. These traditional methods offer straightforward and highly interpretable analytical approaches that help us uncover the physical information embedded in diffusion trajectories.
@@
To address this challenge, machine learning techniques have been introduced to the inference of diffusion parameters due to its powerful ability to handle complex data in recent years. In particular, inferring diffusion parameters essentially constitutes a regression or classification problem, fitting well within the standard supervised learning paradigm. For clarity in our discussion, we have organized these ML approaches into two primary categories in this section: classical machine learning and deep learning. Within classical machine learning, we highlight significant contributions using techniques like feature-based models [117, 118, 119, 120]. Regarding deep learning, we focus on innovative applications that employ deep neural networks, including convolutional neural networks [121, 122, 123, 124, 125, 126, 127, 128, 129], recurrent neural networks [130, 131, 95, 132, 133], graph neural networks [134, 135, 136], to analyze diffusion trajectories and infer the parameters.
@@ 2.1.1 Classical machine learning
Classical machine learning encompasses a range of algorithms and techniques that predate the extensive use of deep neural networks. These methods are grounded in established statistical learning theories and are characterized by their simplicity, interpretability, and lower computational requirements. For the analysis of anomalous diffusion, classical machine learning primarily relies on feature-based models to determine diffusion parameters.
As highlighted in figure 2(a), feature-based models operate on feature vectors transformed from diffusion trajectories and utilize classic machine learning algorithms such as decision trees and multilayer perceptrons for parameter inference. The elements of these trajectory feature vectors are derived from data preprocessing, feature transformation, and feature extraction techniques. By harnessing handcrafted features that capture the essential characteristics of the diffusion process, such as fractal dimension, maximal excursion, and mean gaussianity [72, 73, 74], these models provide a structured way to analyze and interpret diffusion data. Typical related works include the following:
Figure 2: (a) Schematic flowchart of the feature-based methods: a set of features is extracted from raw trajectories and used as the input for feature-based machine learning models, from [75]; (b) A workflow of random forest algorithm for the single trajectory characterization tasks, from [117]; (c) Schematic of the architecture of ELM (top), and corresponding performance in classifying diffusion models (bottom), from [119].
  • A study by G. Muñoz-Gil et al. [117] employs the random forest (RF) algorithm for the single trajectory characterization tasks, which is a classical machine learning method by aggregating results of multiple decision trees through voting or averaging As shown in figure 2(b), upon the supervision of labeled simulated trajectories from various diffusion models, the RF model can be trained to infer diffusion exponent and theoretical model after preprocessing of these trajectories. The trained model is then applied to experimental data, including the movement of mRNA molecules in bacterial cells and membrane receptor diffusion in living cells. The results show that the RF model achieves performance superior to traditional methods, particularly in the analysis of nonergodic diffusion processes.
@@
  • Another application of classical machine learning is the use of extreme learning machine (ELM) in C. Manzo’s work [119]. This method is tailored for single-layer feedforward networks [figure 2(c)], simplifying the learning process by randomly assigning weights between the input and hidden layers. In this study, ELM is integrated with feature engineering to analyze anomalous diffusion trajectories. The input to the ELM model includes features such as time-normalized mean absolute displacement, mean square displacement, and cumulative square displacement. These features are efficiently processed by ELM to infer critical diffusion parameters like diffusion exponent and diffusion model. The model performance underscores the effectiveness of ELM in handling complex, feature-rich datasets, demonstrating its potential to deliver rapid and accurate analysis of diffusion behaviors.
@@
In practice, the effectiveness of these feature-based models heavily depends on the features transformed from trajectory data, such as TA-MSD, displacement moments, or other engineered features that characterize the diffusion dynamics, which will be discussed in detail in section 3.1. As these features typically possess strong statistical or physical significance, they provide feature-based methods with enhanced interpretability compared to deep learning. However, selecting the optimal combination of features is challenging, which requires deep domain expertise and extensive experimental analysis to identify the most significant characteristics of diffusion properties. In this regard, H. Loch-Olszewska et al. underscores the impact of feature choice on machine learning classification of anomalous diffusion [120]. The study shows that while combining multiple feature sets can enhance the performance of classifier, no single feature set can universally excel across all diffusion models. This variability highlights the profound impact of feature selection, and also suggests that feature-based machine learning methods may have limitations in their universality for the analysis of anomalous diffusion.
@@ 2.1.3 Further discussions
While both classic machine learning and deep learning have demonstrated outstanding performance in the inference of diffusion parameters, there are critical considerations that need to be addressed for these methods to reach their full potential. First, the application of machine learning in this field primarily falls within the paradigm of supervised learning, where the task is either a classification or regression problem. Therefore, large and high-quality datasets with accurate labels are required when training these models. However, obtaining labels from experimental data can be challenging. To address this issue, the main current approach is the simulation-based method [117, 118, 119, 120, 137, 122, 123, 124, 121, 125, 128, 126, 127, 130, 131, 95, 132, 133, 134, 136, 135, 138]. This method uses various known mathematical diffusion models, such as CTRW, FBM, and LW, to simulate trajectories. These trajectories can be labeled based on the diffusion models and corresponding simulation parameters. After training the machine learning models on simulated trajectories, these models can be subsequently applied to the analysis of real experimental data. In particular, the first Anomalous Diffusion Challenge [29, 91] specifically tests several competing simulation-based methods on real experimental data [139, 57, 140, 141, 61], like the trajectories of mRNA molecules inside live E. coli cells [139] and single atoms moving in a 1D periodic optical potential [141]. The top-performing methods in this competition show consistent predictions and agree well with experimental values, highlighting that simulation-based approaches can produce robust results across different experimental settings. Details of this competition will be introduced in section 2.3. Although the simulation-based method has shown excellent performance so far, it faces an inherent challenge due to the natural differences between simulated and experimental data. On one hand, experimental trajectories often exhibit higher levels of noise, stochasticity, and complexity compared to their simulated counterparts. On the other hand, many characteristics of experimental trajectories cannot be fully captured and simulated by known diffusion models. This clearly limits the generalizability of machine learning models trained solely on synthetic data. Thus, the improvement and optimization of the simulation-based method will be a critical issue for future research.
@@
    The interpretability of models represents a crucial distinction between classical machine learning and deep learning approaches. Classical techniques like decision trees and random forests offer higher transparency due to their reliance on explicitly defined features. As stated before, these crafted features are selected with substantial domain knowledge. They often root in well-understood physical or statistical properties of the diffusion process, such as TA-MSD and VACF. Therefore, these feature-based representations of anomalous diffusion often prove more robust to differences between experimental and simulation data, due to their more generalized encoding of the trajectory. This level of clarity is invaluable, especially in scientific research where understanding the underlying mechanisms is as crucial as the predictive performance of the model. On the contrary, deep learning models are often criticized for their “black box" nature, as their internal workings remain largely opaque. The lack of interpretability makes it difficult to verify the predictions of deep learning models, sometimes leading to false confident predictions. This is also highlighted in Ref. [129] when dealing with the out-of-distribution dynamics of anomalous diffusion. Nevertheless, it is possible to gain insights into the role of each component within deep neural networks through methods such as ablation studies [142, 143]. These techniques can help elucidate the underlying mechanisms by which neural networks infer diffusion parameters.
@@ 2.2 Segmentation of heterogeneous diffusion dynamics
As highlighted in the introduction, a significant distinction between anomalous diffusion and Brownian motion emerges in complex environments, where a random walker may exhibit varying diffusion states over time due to intrinsic or extrinsic changes, leading to heterogeneous diffusion dynamics [44, 45, 46, 13, 14, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57]. Compared to single-state trajectories with fixed properties, heterogeneous dynamics involve transitions among multiple diffusion states. This makes the accurate analysis of multi-state trajectory data much more complex and thus poses significant challenges in the single trajectory characterization. Segmentation of heterogeneous diffusion dynamics is therefore a crucial task in the analysis of anomalous diffusion. This task involves two primary subtasks:
@@ 2.2.2 Point-wise method
While the sliding window method offers a systematic approach to segmenting trajectories, it inherently suffers from limitations, particularly regarding the predefinition of window length. Considering that the performance of machine learning algorithms heavily depends on the trajectory length [130, 125, 29, 135, 128], the choice of window length is often a trade-off. Longer trajectories provide richer, more comprehensive data, which can be advantageous for capturing overall diffusion patterns. However, using larger windows may smooth out rapid transitions between states, potentially missing important short-term dynamics. Conversely, shorter trajectories tend to lack sufficient information, making it difficult for machine learning models to extract meaningful features. In these cases, smaller windows might increase noise, which can obscure the underlying diffusion process and lead to misclassifications. These constraints become more pronounced when dealing with complex and heterogeneous diffusion dynamics, where transitions between states do not necessarily follow uniform lengths.
@@
Table 1: Overview of selected methods for the segmentation of anomalous diffusion (partially derived from Ref. [152]).
@@
Bayesian deep learning [159] Point-wise ML+Stat. Yes model
@@
Another example of point-wise method is the U-AnDi model proposed by X. Qu et al. [143]. As illustrated in figure 6(c), this model combines dilated causal convolutions to capture long-range dependencies, gated activation units for time-series information filtering, and a U-Net architecture to achieve fine-grained segmentation performance. Similar to the STEP method, U-AnDi predicts diffusion parameters for each point along a trajectory, including the diffusion exponent and diffusion model. As shown in figure 6(d), after training on simulated datasets, U-AnDi is further tested on experimental data involving the diffusion of transmembrane proteins on cell membrane surfaces. These experimental trajectories include two diffusion models: CTRW and FBM, with U-AnDi demonstrating excellent segmentation performance for both. In addition to the STEP and U-AnDi models, other machine learning approaches like RNN [158, 159] and GNN [136] also employ point-wise methods to segment trajectories effectively. Detailed information regarding these models is listed in Table 1.
@@ 2.3 Anomalous Diffusion Challenge
The AnDi Challenge is organized by a group of researchers [164] and is hosted on the CodaLab platform [165]. This competition features two distinct versions. The first iteration in the year 2020 focuses primarily on the inference of diffusion parameters, including the diffusion exponent and diffusion model [29]. While this version includes a segmentation task, its simplistic design does not fully replicate the complexities encountered in real-world scenarios. Consequently, the subsequent iteration of the challenge in the year 2024 shifts its emphasis towards the segmentation of heterogeneous diffusion dynamics [152], aiming to address this complex problem more directly and effectively. Details of the AnDi Challenge are listed in Table 2, including the task descriptions, evaluation metrics, number of teams for each task, and links to related works. The ranking information can be found on the official competition website [91] and is not listed in the table here for brevity. In the following, we provide an overview of each version of the AnDi Challenge, focusing on their specific objectives and respective datasets:
@@
$2^{\rm nd}$ This challenge is organized along two tracks: Video track: For analysis of raw SPT videos. Trajectory track: For the analysis of trajectory data. Each track has two tasks: Task 1 (Single-trajectory task): For each trajectory, participants are supposed to list the changepoints. After that, they need to determine the diffusion coefficient, diffusion exponent, and motion type for each segment. Task 2 (Ensemble task): For each experimental condition, participants are required to provide the model type, number of states, mean and standard deviation of diffusion coefficients, and diffusion exponents for each state. Task 1: JSC, RMSE, MSLE, MAE, F1-Score Task 2: Wasserstein distance Video track: Task 1: 3 teams Task 2: 3 teams Trajectory track: Task 1: 18 teams Task 2: 11 teams [152, 159]
@@
However, despite its significant contributions, the AnDi Challenge highlights a critical gap in this field: the reluctance within the scientific community to share real experimental data openly. Numerous studies on the machine learning characterization of anomalous diffusion rely heavily on the simulation-based method due to the lack of accessible real datasets. This situation constrains the development of machine learning tools that could be effectively applied to actual experimental analysis. To address this issue, there is a pressing need for the SPT research community to actively promote data sharing. In particular, establishing standardized sharing protocols is also important to ensure data quality, consistency, and interoperability. Encouraging the open source of real experimental data under these protocols would not only validate the applicability of these models in practical settings, but also accelerate the discovery of novel insights into the dynamics of anomalous diffusion processes.